• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 08:38:28 on December14,2008

    Entry number 253354

    This entry is a followup to: 253305

    Followups:

    keyword=Compton spectrum is back

    The Compton spectrum (Fig. 1) has been back since the 23:30-01:18 run last night. We again measure an asymmetry of around 7% at the Compton edge.

    Judging by the autologged strip tool from owl shift (Fig. 2), it looks as though the time when the Compton did not appear to be working (no Compton shape to the spectrum, very low asymmetry) coincides with some extremely high rates in the photon detector (blue line). Normal running is a few hundred Hz/uA, but it looks as though the rates were a few thousand Hz/uA for a few hours. This would certainly explain why the characteristic Compton spectrum vanished and why we were measuring such low (i.e. background-diluted) asymmetries.

    What I don't know is why the rates were so high, or what was adjusted to bring them back into the normal range.


    FIGURE 1

    FIGURE 2